# TimeSiam: A Pre-Training Framework for Siamese Time-Series Modeling

### v1

@@ Abstract
Time series pre-training has recently garnered wide attention for its potential to reduce labeling expenses and benefit various downstream tasks. Prior methods are mainly based on pre-training techniques well-acknowledged in vision or language, such as masked modeling and contrastive learning. However, randomly masking time series or calculating series-wise similarity will distort or neglect inherent temporal correlations crucial in time series data. To emphasize temporal correlation modeling, this paper proposes TimeSiam as a simple but effective self-supervised pre-training framework for Time series based on Siamese networks. Concretely, TimeSiam pre-trains Siamese encoders to capture intrinsic temporal correlations between randomly sampled past and current subseries. With a simple data augmentation method (e.g. masking), TimeSiam can benefit from diverse augmented subseries and learn internal time-dependent representations through a past-to-current reconstruction. Moreover, learnable lineage embeddings are also introduced to distinguish temporal distance between sampled series and further foster the learning of diverse temporal correlations. TimeSiam consistently outperforms extensive advanced pre-training baselines, demonstrating superior forecasting and classification capabilities across 13 standard benchmarks in both intra- and cross-domain scenarios.
@@ 1 Introduction
Previous pre-training methods can be roughly categorized into the following two paradigms. As presented in Figure 1, the first one, named masked modeling, enables representation learning by optimizing the model to reconstruct the masked part using the visible context, which has been commonly used in various fields, including natural language processing (Devlin et al., 2018; Raffel et al., 2020) and computer vision (He et al., 2022; Xie et al., 2022; Li et al., 2023a). However, Dong et al. (2023) found that randomly masking a portion of time points will seriously distort vital temporal correlations of time series, making the reconstruction task too difficult to guide representation learning. The other paradigm, contrastive learning, excels in instance-level representation learning, which optimizes the model to identify positive samples from negative ones (Tang et al., 2020; Wickstrøm et al., 2022). A significant criticism of these contrastive approaches is their reliance on careful selection of augmentations to learn useful invariances (Xiao et al., 2020), which is even harder in time series due to the scarcity of widely accepted and effective data augmentation methods (Wen et al., 2020). Also, the instance-level modeling design may fail in capturing fine-grained temporal variations, limiting their practicality to downstream tasks, such as forecasting (Zhang et al., 2022; Wu et al., 2023a).
@@ 2.1 Time Series Self-supervised Pre-training
As a fundamental technique in self-supervised pre-training methods, masked modeling enables deep models to learn essential representations by reconstructing masked parts from the visible context. Drawing inspiration from notable advances in NLP and CV, extensive time series pre-training approaches focus on time series masked modeling, which helps the model learn effective time series representations to facilitate various downstream analysis tasks. For instance, TST (Zerveas et al., 2021) and Ti-MAE (Li et al., 2023b) propose to randomly mask segments and points in time series and pre-train the model with the reconstruction task. PatchTST (Nie et al., 2023) divides time series into several non-overlapping patches and derives a patch-level masked modeling strategy. HiMTM (Zhao et al., 2024) proposes a novel hierarchical masked time series pre-training method to capture the multi-scale nature of time series. Additionally, SimMTM (Dong et al., 2023) introduces a multi-masking modeling paradigm, which reconstructs original time series through the weighted aggregation of multiple masked time series, thus being able to learn both points-wise and series-wise temporal representations. Despite these advances, all of these approaches solely focus on the modeling of one individual time series, disregarding the intrinsic temporal correlations and dynamical variations of the whole time series. In contrast, TimeSiam proposes to reconstruct the current sub-series based on past observations, which can naturally integrate the time-dependent information during reconstruction pre-training.
@@
Unlike masked modeling, this approach enables model pre-training by optimizing the similarity among instance-level representations. It leverages different data augmentations to construct positive and negative pairs from data, where positive pairs are optimized to be close to each other and negative pairs are encouraged to be distant from each other during pre-training (Tang et al., 2020; He et al., 2020; Chen & He, 2021; Gao et al., 2021). Current time series contrastive learning methods are mainly based on diverse data augmentations tailored to the domain-specific characteristics of time series. TS2Vec (Yue et al., 2022) divides time series into patches, defining contrastive tasks at both the individual instance and patch levels. Mixing-up (Wickstrøm et al., 2022) exploits a data augmentation scheme in which new samples are generated by mixing two data samples. LaST (Wang et al., 2022) aims to separate seasonal and trend components in time series data within the latent space. Additionally, CoST (Woo et al., 2022) utilizes contrastive losses in both time and frequency domains to learn distinct seasonal and trend representations. Furthermore, TF-C (Zhang et al., 2022) introduces a novel time-frequency consistency architecture, optimizing for proximity between time-based and frequency-based representations of the same data sample. However, existing contrastive learning methods for time series heavily rely on intricate data augmentation techniques to generate diverse views of the original data for self-supervision. Also, the instance-level representation learning may fall short in downstream low-level tasks. In TimeSiam, we utilize the native temporally distanced subseries to build reconstruction tasks, thereby freeing from complex augmentation techniques and also considering the low-level representation.
@@ 2.2 Siamese Networks
Siamese networks (Bromley et al., 1993) are particular neural network architectures with shared model parameters. This design makes Siamese networks well-suited for comparing and distinguishing two input samples and has been widely used in contrastive learning to model the relationship between paired samples (Chen & He, 2021). The combination of Siamese networks and contrastive learning has been widely used in many applications, particularly in tasks requiring instance-level representations (Chen et al., 2020; He et al., 2020; Wang et al., 2023). However, in the field of time series pre-training, this combination generally focuses on recognizing subtle differences between various augmented views of the series itself, overlooking the essence of time series, that is temporal correlation modeling. In this paper, we explore using shared-weight Siamese autoencoders to establish correlations between past and current subseries. This methodology enables a more efficient understanding of temporal relations in time series and enforces the model to learn time-dependent representations.
@@ 3.1 Pre-training
After constructing mask-augmented Siamese pairs, as shown in Figure 2, we further integrate learnable lineage embeddings during pre-training to effectively capture the disparity among different Siamese pairs. This design can enhance the model’s capacity to extract diverse temporal-related representations. Given $N$ learnable lineage embeddings $\{\mathbf{e}^{\text{lineage}}_{i}\}_{i=1}^{N},\mathbf{e}^{\text{lineage}}_{i}\in\mathbb{R}^{1\times D}$, for the past sample $\mathbf{{x}}^{\text{past}}$, we apply the $\operatorname{LineageMatching}(\cdot)$ function to dynamically match a certain lineage embedding based on its temporal distance $d$ to the current series. As for the current sample $\mathbf{\widetilde{x}}^{\text{curr}}$, we use a special lineage embedding to represent a degeneration situation as $d=0$:
@@
where $\mathbf{e}^{\text{lineage}}_{0}\in\mathbb{R}^{1\times D}$ is the specific embedding for current subseries and $\mathbf{z}^{\text{past}},\mathbf{\widetilde{z}}^{\text{curr}}\in\mathbb{R}^{T\times D}$ denote the embedded Siamese features. Note that lineage embeddings are used to identify the temporal distance between Siamese subseries. It is shared along the time dimension when being added to Siamese subseries features. Here, $\oplus$ represents the addition operation with the temporal dimension broadcast.
@@ 3.2 Fine-tuning
$\begin{split}\overline{\mathbf{h}}_{e}&=\operatorname{Concat}\left(\mathbf{h}_{e,0},\mathbf{h}_{e,1},\ldots,\mathbf{h}_{e,k}\right),\\ \text{where}\ \mathbf{h}_{e,i}&=\operatorname{Encoder}\left(\operatorname{Embed}(\mathbf{x}_{i})\oplus\mathbf{e}^{\text{lineage}}_{{\operatorname{\text{\tiny{LineageMatching}}}(\tiny{iT}})}\right).\end{split}$ (8)
@@ 4.1 Experimental Setup
Backbone
We use the advanced time series models across various tasks as Siamese encoders to evaluate the efficacy of our pre-training methods. In particular, we utilized iTransformer (Liu et al., 2024), and PatchTST (Nie et al., 2023) as the encoder for time series forecasting. TCN (Bai et al., 2018) is used as the backbone for the classification task (Wang et al., 2023). Note that to ensure a fair comparison, we unify the encoder backbone of our model and all the baselines. The results with a unified encoder generally surpass the results reported by themselves across all baselines.
Baselines
We compare our TimeSiam with eight advanced self-supervised time series pre-training baselines under the in-domain setting, including contrastive learning methods: COMET (2023), TF-C (2022), LaST (2022), CoST (2022), TS2Vec (2022), and masked modeling methods: SimMTM (2023), Ti-MAE (2023b), TST (2021). In the cross-domain setting, due to the large-scale dataset TSLD will bring huge experiment costs, we only select part of the baselines (the efficient ones) for comparisons. Besides, COMET (2023) is specifically designed for medical time series and we also exclude it from cross-domain evaluation.
@@ 4.2 Main Results
In Figure 3, we summarize the performance of our TimeSiam in both in- and cross-domain scenarios for forecasting ($x$-axis) and classification tasks ($y$-axis). For each scenario, TimeSiam exhibits significant improvement over other established strong self-supervised baselines for time series.
@@ 4.3 Forecasting
We investigate the effectiveness of TimeSiam by integrating it with state-of-the-art time series forecasting models: PatchTST (Nie et al., 2023) and iTransformer (Liu et al., 2024). As shown in Table 2, TimeSiam can further boost the performance of these models, achieving an average MSE reduction of 5.7% and 3.2%, even when these advanced base models already have excellent forecasting capabilities Random init. Significantly, the masked modeling-based approach generally outperforms the contrastive-based approach in the forecasting task. However, our TimeSiam still performs the best prediction capability among all existing state-of-the-art self-supervised baseline methods.
@@
(a) Siamese Sampling (b) Subseries Reconstruction
Sampling range $(\max d)$
Lineage types $N$ Masking rules Masked ratio
$w/o$ 0.462 $w/o$ 0.457 binomial 0.459 $w/o$ 0.460
1 0.454 2 0.456 continuous 0.459 15% 0.459
3 0.454 3 0.453 mask last 0.461 25% 0.453
6 0.453 6 0.455 binomial (c) 0.457 50% 0.455
12 0.455 continuous (c) 0.453 75% 0.457
@@ 4.4 Classification
To further investigate the generalizability of the representations learned by TimeSiam, we examined the impact of in-domain pre-training on classification tasks within the medical domain, following the setup in (Wang et al., 2023). Results in Table 5 demonstrate competitive outcomes achieved by both COMET (Wang et al., 2023) and SimMTM (Dong et al., 2023). This can be attributed to the elaborative designs of their approaches, where COMET incorporates domain-specific knowledge into its design and SimMTM models both high-level and low-level representations. Compared with these competitive baselines, our TimeSiam, characterized by its simplicity and generality, consistently achieves remarkable results.
@@ 4.5 Ablation Studies
We explore the key hyper-parameters of Siamese modeling: the maximum sampling distance between the past and current Siamese subseries (Sampling range $\max d$) and the size of the lineage embedding set (Lineage types $N$). Table 4(a) demonstrates that integrating past-to-current Siamese modeling outperforms self-reconstruction modeling in time series pre-training. Furthermore, expanding the sampling range of Siamese subseries reasonably significantly enhances performance, underscoring the critical role of Siamese modeling in achieving optimal fine-tuning results. Also, employing lineage embeddings can further boost the model’s performance.
@@
Figure 4: Fine-tuning the pre-trained model to the inputs with extended length, where the input length is selected from {96, 192, 288, 384, 576}. The MSE averaged from 4 forecasting horizons {96, 192, 336, 720} is reported. Additional results can be found in Appendix F.
4.6 Analysis Experiment
Data Scale and Model Capacity

### v2

@@ Abstract
Time series pre-training has recently garnered wide attention for its potential to reduce labeling expenses and benefit various downstream tasks. Prior methods are mainly based on pre-training techniques well-acknowledged in vision or language, such as masked modeling and contrastive learning. However, randomly masking time series or calculating series-wise similarity will distort or neglect inherent temporal correlations crucial in time series data. To emphasize temporal correlation modeling, this paper proposes TimeSiam as a simple but effective self-supervised pre-training framework for Time series based on Siamese networks. Concretely, TimeSiam pre-trains Siamese encoders to capture intrinsic temporal correlations between randomly sampled past and current subseries. With a simple data augmentation method (e.g. masking), TimeSiam can benefit from diverse augmented subseries and learn internal time-dependent representations through a past-to-current reconstruction. Moreover, learnable lineage embeddings are also introduced to distinguish temporal distance between sampled series and further foster the learning of diverse temporal correlations. TimeSiam consistently outperforms extensive advanced pre-training baselines, demonstrating superior forecasting and classification capabilities across 13 standard benchmarks in both intra- and cross-domain scenarios. Code is available at https://github.com/thuml/TimeSiam.
Machine Learning, ICML
@@ 1 Introduction
Previous pre-training methods can be roughly categorized into the following two paradigms. As presented in Figure 1, the first one, named masked modeling, enables representation learning by optimizing the model to reconstruct the masked part using the visible context, which has been commonly used in natural language processing (Devlin et al., 2018; Raffel et al., 2020) and computer vision (He et al., 2022; Xie et al., 2022; Li et al., 2023a). However, Dong et al. (2023) found that randomly masking a portion of time points will seriously distort vital temporal correlations of time series, making the reconstruction task too difficult to guide representation learning. The other paradigm, contrastive learning, excels in instance-level representation learning, which optimizes the model to identify positive samples from negative ones (Tang et al., 2020). A significant criticism of these contrastive approaches is their reliance on careful augmentations selection to learn useful invariances (Xiao et al., 2021), which is even harder in time series due to the scarcity of widely accepted and effective data augmentation methods (Wen et al., 2020). Also, the instance-level modeling design may fail in capturing fine-grained temporal variations, limiting their practicality to downstream tasks.
@@ 2.1 Time Series Self-supervised Pre-training
As a fundamental technique in self-supervised pre-training methods, masked modeling enables deep models to learn essential representations by reconstructing masked parts from the visible context. Drawing inspiration from notable advances in NLP and CV, extensive time series pre-training approaches focus on time series masked modeling, which helps the model learn effective time series representations to facilitate various downstream analysis tasks. For instance, TST (Zerveas et al., 2021) and Ti-MAE (Li et al., 2023b) propose to randomly mask segments and points in time series and pre-train the model with the reconstruction task. PatchTST (Nie et al., 2023) divides the temporal dimension into multiple patches and treats time series as independent variates. Additionally, it incorporates a non-overlapping patch-level masked self-supervised strategy for temporal representation learning. HiMTM (Zhao et al., 2024) proposes a novel hierarchical masked time series pre-training method to capture the multi-scale nature of time series. Additionally, SimMTM (Dong et al., 2023) introduces a multi-masking modeling paradigm, which reconstructs original time series through the weighted aggregation of multiple masked time series, thus being able to learn both points-wise and series-wise temporal representations. Despite these advances, all of these approaches solely focus on the modeling of one individual time series, disregarding the intrinsic temporal correlations and dynamical variations of the whole time series. In contrast, TimeSiam proposes to reconstruct the current sub-series based on past observations, which can naturally integrate the time-dependent information during reconstruction pre-training.
@@
Unlike masked modeling, this approach enables model pre-training by optimizing the similarity among instance-level representations. It leverages different data augmentations to construct positive and negative pairs from data, where positive pairs are optimized to be close to each other and negative pairs are encouraged to be distant from each other during pre-training (Tang et al., 2020; He et al., 2020; Chen & He, 2021; Gao et al., 2021). Current time series contrastive learning methods are mainly based on diverse data augmentations tailored to the domain-specific characteristics of time series. CPC (Oord et al., 2018) introduced contrastive predictive coding, which uses model-predicted timesteps as positive samples and randomly-sampled timesteps as negative samples to obtain advantageous time series representations for downstream tasks. Franceschi et al. (2019) combined a causal dilated convolutions-based encoder with a novel triplet loss that employs time-based negative sampling. TNC (Tonekaboni et al., 2021) learns the representations by ensuring that signals from within a neighborhood are distinguishable from the distribution of non-neighboring signals in the latent space using a debiased contrastive loss. TS2Vec (Yue et al., 2022) divides time series into patches, defining contrastive tasks at both the individual instance and patch levels. Mixing-up (Wickstrøm et al., 2022) exploits a data augmentation scheme in which new samples are generated by mixing two data samples. LaST (Wang et al., 2022) aims to separate seasonal and trend components in time series data within the latent space. Additionally, CoST (Woo et al., 2022) utilizes contrastive losses in both time and frequency domains to learn distinct seasonal and trend representations. Furthermore, TF-C (Zhang et al., 2022) introduces a novel time-frequency consistency architecture, optimizing for proximity between time-based and frequency-based representations of the same data sample. However, existing contrastive learning methods for time series heavily rely on intricate data augmentation techniques to generate diverse views of the original data for self-supervision. Also, the instance-level representation learning may fall short in downstream low-level tasks. In TimeSiam, we utilize the native temporally distanced subseries to build reconstruction tasks, thereby freeing from complex augmentation techniques and also considering the low-level representation.
@@ 2.2 Siamese Networks
Siamese networks (Bromley et al., 1993) are particular neural network architectures with shared model parameters. This design makes Siamese networks well-suited for comparing and distinguishing two input samples based on a single neural network. They have been widely used in contrastive learning to model the relationship between paired samples (Chen & He, 2021). The combination of Siamese networks and contrastive learning has been widely used in many applications, particularly in tasks requiring instance-level representations (Chen et al., 2020; He et al., 2020; Wang et al., 2023). However, in the field of time series pre-training, this combination generally focuses on recognizing subtle differences between various augmented views of the series itself, overlooking the essence of time series, that is temporal correlation modeling. In this paper, we explore using shared-weight Siamese autoencoders to establish correlations between past and current subseries. This methodology enables a more efficient understanding of temporal relations in time series and enforces the model to learn time-dependent representations.
@@ 3.1 Pre-training
After constructing mask-augmented Siamese pairs, as shown in Figure 2, we further integrate learnable lineage embeddings during pre-training to effectively capture the disparity among different Siamese pairs. This design can enhance the model’s capacity to extract diverse temporal-related representations. Given $N$ learnable lineage embeddings $\{\mathbf{e}^{\text{lineage}}_{i}\}_{i=1}^{N},\mathbf{e}^{\text{lineage}}_{i}\in\mathbb{R}^{1\times D}$ and $D$ represents the dimension of lineage embeddings. For the past sample $\mathbf{{x}}^{\text{past}}$, we apply the $\operatorname{LineageMatching}(\cdot)$ function to dynamically match a certain lineage embedding based on its temporal distance $d$ to the current series. As for the current sample $\mathbf{\widetilde{x}}^{\text{curr}}$, we use a special lineage embedding to represent a degeneration situation as $d=0$:
@@
where $\mathbf{e}^{\text{lineage}}_{0}\in\mathbb{R}^{1\times D}$ is the specific embedding for current subseries and $\mathbf{z}^{\text{past}},\mathbf{\widetilde{z}}^{\text{curr}}\in\mathbb{R}^{T\times D}$ denote the embedded Siamese features. Different base models correspond to different $\operatorname{Embed}(\cdot)$. Regarding PatchTST (Nie et al., 2023), the patch-wise embedding function $\operatorname{PatchEmbed}(\cdot)$ is used to divide each variable into several patches and each patch is mapped to a patch token. As for iTransformer (Liu et al., 2024), it uses the variable-wise embedding $\operatorname{VariateEmbed}(\cdot)$ and maps the entire variable into a temporal token. Note that lineage embeddings are used to identify the temporal distance between Siamese subseries. It is shared along the time dimension when being added to Siamese subseries features. Here, $\oplus$ represents the addition operation with the temporal dimension broadcast.
@@ 3.2 Fine-tuning
$\begin{split}\overline{\mathbf{h}}_{e}&=\operatorname{Concat}\left(\mathbf{h}_{e,0},\mathbf{h}_{e,1},\ldots,\mathbf{h}_{e,k}\right),\\ \text{where}\ \mathbf{h}_{e,i}&=\operatorname{Encoder}\left(\operatorname{Embed}(\mathbf{x}_{i})\oplus\mathbf{e}^{\text{lineage}}_{{\operatorname{\text{\tiny{LineageMatching}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tiny{iT}}})}}\right).\end{split}$ (8)
@@ 4.1 Experimental Setup
Backbone
We use the advanced time series models across various tasks as Siamese encoders to evaluate the efficacy of our pre-training methods. In particular, we utilized iTransformer (Liu et al., 2024) and PatchTST (Nie et al., 2023) as the encoder for time series forecasting following their original configurations. The patch length and stride were both set to 12 without any overlap. TCN (Bai et al., 2018) is used as the backbone for the classification task (Wang et al., 2023). Note that to ensure a fair comparison, we unify the encoder backbone of our model and all the baselines. The results with a unified encoder generally surpass the results reported by themselves across all baselines.
Baselines
We compare our TimeSiam with eight advanced self-supervised time series pre-training baselines under the in-domain setting, including contrastive learning methods: COMET (2023), TF-C (2022), LaST (2022), CoST (2022), TS2Vec (2022), TNC (2021), CPC (2018) and masked modeling methods: SimMTM (2023), Ti-MAE (2023b), TST (2021), and a patch-wise self-surpervised masked modeling method, PatchTST, proposed by (Nie et al., 2023).
It should be noted that some baselines such as CPC (Oord et al., 2018), TNC (Tonekaboni et al., 2021), etc. are not validated in the iTransformer backbone because their specific design based on internal modeling of series is not applicable to iTransformer (Liu et al., 2024), which models the entire sequence as a whole temporal token and focuses on modeling relationships between different variables. In the cross-domain setting, due to the large-scale dataset TSLD will bring huge experiment costs, we only select part of the baselines (the efficient ones) for comparisons. Besides, COMET (2023) is specifically designed for medical time series and we also exclude it from cross-domain evaluation. More implementation details can be found in Appendix A.
@@ 4.2 Main Results
As shown in Figure 3,we summarize the performance of our TimeSiam in both in- and cross-domain scenarios for two mainstream time series analysis tasks: time series forecasting ($x$-axis) and classification ($y$-axis). For each scenario, TimeSiam exhibits significant improvement over other established strong self-supervised baselines for time series.
@@ 4.3 Forecasting
We investigate the effectiveness of TimeSiam by integrating it with state-of-the-art time series forecasting models: PatchTST (Nie et al., 2023) and iTransformer (Liu et al., 2024). As shown in Table 2, TimeSiam can further enhance model performance, achieving an average MSE reduction of 5.7% and 2.5% across all forecasting benchmarks, even though these advanced models already exhibit excellent forecasting capabilities from random initialization. Significantly, it is evident that the masked modeling-based approach overall outperforms the contrastive-based approach in the forecasting task. This can be attributed to the advantages gained through series reconstruction to learn low-level temporal representation. However, our TimeSiam still performs the best forecasting capability among all existing state-of-the-art self-supervised baseline methods.
@@
      (a) Siamese Sampling       (b) Subseries Reconstruction
Sampling range $(\max d)$
      Lineage types $N$       Masking rules       Masked ratio
      $w/o$       0.462       $w/o$       0.457       binomial       0.459       $w/o$       0.460
      1       0.454       2       0.456       continuous       0.459       15%       0.459
      3       0.454       3       0.453       mask last       0.461       25%       0.453
      6       0.453       6       0.455       binomial (c)       0.457       50%       0.455
      12       0.455       continuous (c)       0.453       75%       0.457
@@ 4.4 Classification
To further investigate the generalizability of the representations learned by TimeSiam, we examined the impact of in-domain pre-training on classification tasks within the medical domain, following the setup in (Wang et al., 2023). Results in Table 4 demonstrate competitive outcomes achieved by both COMET (Wang et al., 2023) and SimMTM (Dong et al., 2023). This can be attributed to the elaborative designs of their approaches, where COMET incorporates domain-specific knowledge into its design and SimMTM models both high-level and low-level temporal representations. Compared with these competitive baselines, our TimeSiam, characterized by its simplicity and generality, consistently achieves remarkable results. In all classification benchmarks, our proposed TimeSiam consistently enhances the average classification accuracy by 11.5% compared to random initialization, surpassing other baseline methods.
@@ 4.5 Ablation Studies
We explore the key hyper-parameters of Siamese modeling: the maximum sampling distance between the past and current Siamese subseries (Sampling range $\max d$) and the size of the lineage embedding set (Lineage types $N$). The maximum sampling range is determined by the length of the subseries ($T$) and a hyperparameter ($r$), resulting in $\max d=T\times r$. Table 6(a) demonstrates that integrating past-to-current Siamese modeling outperforms self-reconstruction modeling in time series pre-training. Furthermore, expanding the sampling range of Siamese subseries reasonably significantly enhances performance, underscoring the critical role of Siamese modeling in achieving optimal fine-tuning results.
@@
Lineage embeddings
As shown in Table 8, we can observe that lineage embeddings play a vital role in enhancing forecasting performance. This comes from the inherent ability of lineage embeddings to distinguish temporal distances between subseries during the pre-training phase. As a result, they facilitate the learning of diverse temporal correlations. Consequently, the inclusion of lineage embeddings has been shown to be effective in improving performance when fine-tuning models for a variety of downstream tasks.
Figure 4: Fine-tuning the pre-trained model to the inputs with extended length {96, 192, 288, 384, 576} based on iTransformer (Liu et al., 2024). The MSE averaged from all predicted horizons $\{96,192,336,720\}$ is reported. Additional results are in the Appendix F.
@@
Table 8: Ablations on lineage embeddings. The MSE averaged from 4 forecasting horizons {96, 192, 336, 720} is reported here.
Datasets Random init. TimeSiam w/o Lineage
ETTh1 0.473 0.429 0.433 $\downarrow$
ETTm1 0.390 0.374 0.378 $\downarrow$
Weather 0.259 0.252 0.256 $\downarrow$
Traffic 0.490 0.453 0.457 $\downarrow$
Exchange 0.36 0.353 0.365 $\downarrow$
4.6 Analysis Experiment
Data Scale and Model Capacity